Solve each equation

66. (x – 3)2 + (x + 2)2 = 17

Answer by je02(2)

(x – 3)2 + (x + 2)2 = 17
2x - 6 + 2x + 4 = 17
4x - 2 = 17
4x = 17+2
4x = 19
divide both sides by 4
x = 19/4
to check:
(19/4 - 3)2 + (19/4 + 2)2 = 17
(19/4 - 12/4)2 + (19/4 + 8/4)2 = 17
(7/4)2 + (27/4)2 = 17
(14/4) + 54/4 = 17
68/4 = 17
17 = 17